# Gaussian Splatting Lucas-Kanade

### v1

@@ 4.1 Forward warp field velocity
With the aforementioned problem statement, we show that the velocity field can be derived from the forward warp field. Restating the objective, we seek to find a warp field $\mathcal{W}$ that transforms the canonical Gaussians $\mathcal{G}$ forward in time such that it minimizes the appearance loss and scene flow loss at the target time step. To do this, we derive the velocity of the forward warp field as analogous to the twist Jacobians. The velocity field $v(\mathcal{G};t)$ expresses the changes in Gaussian parameters in the world coordinate at a given time $t$. Gaussian Splatting maintains an explicit canonical Gaussian space so intuitively the velocities can be computed from its forward warp field $\mathcal{W}_{c\rightarrow}(\mathcal{G},t)$ as:
$v(\mathcal{G};t)=\frac{\partial\mathcal{W}_{c\rightarrow}(\mathcal{G},t)}{\partial t}\Big{|}_{(\mathcal{W}_{c\leftarrow}(\mathcal{G},t);t)}.$ (9)
The explicit expression of 3D Gaussians ensures that the scene flow mapping between the adjacent frames is bijective. In other words, the resulting correspondences among the 3D points in each frame are cyclic consistent since they are deformed from the same canonical Gaussian. This allows us to analytically compute the velocity field $v(\mathcal{G};t)$ for each Gaussian in the canonical space.
The velocity field $v(\mathcal{G};t)$ requires an analytical representation of the warp field’s gradient concerning time and warp parameters. Although the direct analytical solution is costly to compute, the chain rule can be applied to decompose the compute into two partial terms: the feature gradient of the warp field with respect to the corresponding Gaussian parameter, and the analytical warp Jacobian Aoki2019Pointnetlk:Pointnet ; Wang_2023_CVPR :
$v(\mathcal{G};t)=\boldsymbol{J}^{-1}_{\mathcal{G}}\frac{\partial\mathcal{W}_{c\rightarrow}(\mathcal{G};t)}{\partial t}=\frac{\partial\mathcal{W}_{c\rightarrow}(\mathcal{G};t)}{\partial\mathcal{G}}\frac{\partial\mathcal{W}_{c\rightarrow}(\mathcal{G};t)}{\partial t}.$ (10)
The warp field feature gradient $\frac{\partial\mathcal{W}_{c\rightarrow}(\mathcal{G};t)}{\partial t}$ describes how each output Gaussian parameter offsets changes with respect to the input canonical Gaussian parameters. The warp Jacobian $\frac{\partial\mathcal{W}_{c\rightarrow}(\mathcal{G};t)}{\partial\mathcal{G}}$ describes how changes in the warp field output affect the transformation on the Gaussians.
We note that the expression holds for any Gaussian parameter that is an input to the deformation network $\mathcal{W}_{c\rightarrow}$ as long as the respective input and output dimensions are the same. However, in practice, the warp Jacobian can experience numerical instabilities with $det(\boldsymbol{J}_{\mathcal{G}})<\epsilon$, especially when the canonical Gaussian scene has not stabilized. However, discarding unstable motions also affects the downstream densification and pruning process. We instead replace $\boldsymbol{J}^{-1}_{\mathcal{G}}$ with Moore–Penrose pseudoinverse operation $\boldsymbol{J}^{+}_{\mathcal{G}}$. The scaling terms and spherical harmonics of Gaussians have minimal changes between adjacent frames, we thus choose to exclude them from the compute for efficiency. Overall, we compute the velocity fields for Gaussians’ displacement and rotation as:
$v(\boldsymbol{p};t)=\boldsymbol{J}^{+}_{\boldsymbol{p}}\frac{\partial\mathcal{W}_{c\rightarrow}(\boldsymbol{p};t)}{\partial t};v(\boldsymbol{q};t)=\boldsymbol{J}^{+}_{\boldsymbol{q}}\frac{\partial\mathcal{W}_{c\rightarrow}(\boldsymbol{q};t)}{\partial t},$ (11)
@@
where $v(\boldsymbol{p};t):\mathbb{R}^{3}\times\mathbb{R}\rightarrow\mathbb{R}^{3}$ and $v(\boldsymbol{q};t):\mathbb{R}^{4}\times\mathbb{R}\rightarrow\mathbb{R}^{4}$ respectively describes the positional and rotational velocity of a Gaussian in the world coordinate.
@@ 4.2 Scene field from time integration
Time integration on velocity field. Given the velocity field from equation 10, and 3D Gaussians $\mathcal{G}(t)$ observed at time step $t$, we can apply time integration using a Runge-Kutta Batschelet1952UberDifferentialgleichungen numerical solver on the velocity field to obtain the offset parameters for Gaussians at time $(t+\Delta t)$:
@@
We then train the warp field network to minimize the difference between the predicted scene flow derived from the analytical velocity field, and the ground truth scene flow decomposed into optical flow and depth describing the in-plane and out-of-plane Gaussian motions.
@@
For optical flow rendering, we want to calculate the motion’s influence on pixel shifts from all Gaussians in the world coordinate. In the original 3D Gaussian Splatting, a pixel’s color is the weighted sum of the projected 2D Gaussian’s radiance contribution. Analogous to this formulation, the optical flow at a queried pixel is the weighted sum of the 2D Gaussians’ contribution to the pixel shift:
$\hat{flow}^{G}_{t,t+\Delta t}=\Sigma_{i\in N}flow_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})={\sum}_{i=1}w_{i}[\Sigma_{i,t_{2}}\Sigma^{-1}_{i,t_{1}}(x_{t_{1}}-\mu_{i,t_{1}})+\mu_{i,t_{2}}-x_{t_{1}}],$ (13)
where $flow_{i}$ denotes the optical flow of Gaussian $\mathcal{G}_{i}$, and $w_{i}=\frac{T_{i}\alpha_{i}}{\Sigma_{i}T_{i}\alpha_{i}}$ denotes the weighing factor of each Gaussian from $\alpha$-blending.
Similarly, the per-pixel z-depth estimates $\hat{d}$ can be rendered from the discrete volume rendering approximation accounting for the per-Gaussian contributions:
$\hat{D}^{G}_{t,t+\Delta t}=\Sigma_{i\in N}d_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (14)
where $d_{i}$ denotes the z-depth value of Gaussian $\mathcal{G}_{i}$ from the viewing space. We normalize the optical flow and depth estimates for numerical stability. The optical flow and depth maps are rasterized simultaneously in the same forward pass with color.
@@ 5 Experiments
In this section, we first provide the implementation details of the proposed warp field regulation and then validate our proposed method on four dynamic scene datasets captured with different levels of camera movements. We showcase the quantitative and qualitative results of the Dynamic Scenes dataset Yoon2020NovelCamera , Planoptic Video dataset Li_2021_CVPR , HyperNeRF dataset Park2021HyperNeRF:Fields , and DAVIS dataset Perazzi2016ASegmentation . Our method achieves state-of-the-art results on both static and dynamic cameras.
@@ 5.1 Implementation details
We implement the warp field using PyTorch Paszke2019PyTorch:Library , leveraging its Autodiff library for gradient and Jacobian computations. The framework is optimized with Adam Kingma2015Adam:Optimization as with 3DGS kerbl20233d on a NVIDIA A100 Tensor Core NvidiaCorporation2024NVIDIACore . We use the Torchdiffeq torchdiffeq library for numerical integration. Modified from DeformableGS, the inputs to the warp field include 3D Gaussian positions ($\boldsymbol{p}$), time ($t$), and 3D Gaussian rotations ($\boldsymbol{q}$) in quaternions. Velocity derivations are enabled by removing the stop gradient operation from the network inputs. Positional encoding mildenhallnerf is applied to extend inputs’ frequency band zheng2022trading . We refer readers to Appendix A.2 for more implementation details.
@@ 5.2 Evaluation
Benchmarked datasets. We evaluate our method using real-world monocular datasets with varying camera motion. These include Dynamic Scene Yoon2020NovelCamera , captured by a stationary multi-view 8 camera rig with significant scene motion; Plenoptic Video Li_2021_CVPR , captured using a static rig with 21 GoPro cameras GoPro2023Go7 ; Hypernerf Park2021HyperNeRF:Fields , which captures objects with moving topologies from a moving camera, suitable for quasi-static reconstruction; and sequences from DAVIS 2017 Perazzi2016ASegmentation dataset containing near-static monocular videos.
Quantitative baseline comparisons. We evaluate our method’s novel-view synthesis capabilities on three datasets: Plenoptic Videos, Hypernerf, and Dynamic Scenes, ordered by decreasing Effective Multi-view Factor (EMF) NEURIPS2022_emfs . These datasets present increasing challenges, with Dynamic Scenes posing the greatest difficulty due to its low EMF and unreliable object motion. We assess visual quality using LPIPS zhang2018perceptual , SSIM wang2004image , and PSNR Welstead2009FractalTechniques , comparing our method to static Gaussian Splatting kerbl20233d , DeformableGS Yang2023DeformableReconstruction , and 4DGaussians wu20234dgaussians , as shown in Table 4 and Table 2. Our method consistently outperforms these baselines on scenes with higher EMFs (Plenoptic Videos and Hypernerf).
@@
To further assess our method’s performance, we compare its novel-view synthesis results with NSFF Li2021NeuralScenes , NR-NeRF tretschk2021non , Nerfies park2021nerfies , and Flow-supervised NeRF Wang_2023_CVPR on the Dynamic Scenes dataset, see Table 1. These comparisons highlight the effectiveness of our analytical scene flow approach compared to other flow-guided rendering frameworks. Notably, our method outperforms NSFF, which uses a similar supervision style but a different scene flow derivation. The comparison with Flow-supervised NeRF, which also uses analytical scene flow derived from an implicit radiance field, serves as an ablation study, demonstrating the effectiveness of our derived scene flow on an explicit scene representation. Finally, we support the aforementioned decomposed structural supervision as in Table 4. More discussions on alternative setups are provided in Appendix Sec. A.3.
@@
Table 4: Ablation study on scene flow decomposition on "Playground" scene from Dynamic Scenes dataset. See Sec. 5.2 for detailed descriptions of each ablated design.
@@
Trajectories from scene flow fields. We visualize the scene flow trajectories from Dynamic Scenes Dataset sequences to assess the quality and smoothness of the derived flow field. The Gaussians are subsampled from the dynamic regions of the canonical Gaussian space, and then time integrated to produce their respective displacements across time, visualized as colored trajectories in Fig. 2. We note that the trajectories are smooth and follow the expected motions of the dynamic objects. The visualized trajectories of the sequences demonstrate the framework’s potential to be extended for tracking in dynamic 3D Gaussian scenes. More trajectories on the DAVIS dataset will be shown in the appendix.
Geometric consistency. Fig. 4 compares our method with DeformableGS, which learns Gaussians’ time dependency without scene flow regularization. While DeformableGS renders visually accurate novel views, the underlying scene structures are inaccurate, leading to blurry dynamic objects and artifacts in depth maps. Our method, leveraging analytical scene flow regularization, achieves more accurate geometries, reflected in both visual and quantitative results (Table 1). This also enables accurate Gaussian motion tracking (Fig. 2). To explore alternative design choices, we compare with direct flow and depth supervision on DeformableGS (Fig. 1). These comparisons indicate that our approach, which enforces regularization on the expected scene flow over a continuous time frame, results in more plausible scene structures compared to discrete time step supervision.
Figure 5: Our method struggles with insufficient point cloud initialization due to reliance on non-rigid warping of scene geometry. In the ’skating’ scene, this results in unstable geometry at certain angles. Inaccurate camera calibrations also degrade our method, as shown in the ’toby-sit’ sequence, where miscalibration distorts the scene geometry.
@@ A.1 Scene Flow Regularization Formulation
Recall that in addition to the photometric regularization as detailed in the original 3D Gaussian Splatting kerbl20233d , $\mathcal{L}_{scene}$ is used as the structural regularization for the derived scene flow consisting of weighted depth and optical flow terms $\mathcal{L}_{scene}=\alpha\mathcal{L}_{o_{t\rightarrow t+1}}+\beta\mathcal{L}_{d_{t\rightarrow t+1}}$.
The optical flow term minimizes the absolute error between the rendered optical flow as computed from accounting for all Gaussians’ 2D projected motion contributions to each rendered pixel, and the reference flow from RAFT Teed2020RAFTRA . In particular, the optical flow term is expressed as:
$\mathcal{L}_{o_{t\rightarrow t+1}}=\frac{1}{n}\sum_{j=1}^{n}|o_{j}-\hat{o}_{j}|,$ (15)
where $j$ denotes the queried pixel from a rasterized image containing $n$ total pixels.
As noted in SparseNeRF wang2023sparsenerf , monocular depth maps are scale-ambiguous thus directly using monocular depth priors with $p$-norm regularizations hinders the spatial coherency. We apply a local depth ranking loss on the rendered depth maps. We transfer the depth ranking knowledge from monocular depth $d_{mono}$ to rendered $d_{r}$ with:
$\mathcal{L}_{d_{t\rightarrow t+1}}=\sum_{d^{k_{1}}_{mono}<d^{k_{2}}_{mono}}max(d^{k_{1}}_{r}-d^{k_{2}}_{r}+m,0),$ (16)

### v2

@@ 4.1 Forward warp field velocity
To address the challenge of transforming canonical Gaussians $\mathcal{G}$ over time, our goal is to derive the velocity field from the forward warp field. Specifically, we seek to identify a warp field $\mathcal{W}$ that advances the canonical Gaussians in time, with an underlying velocity field $v(\mathcal{G};t)$ that quantifies the changes in Gaussian parameters within the world coordinate system at a specific time $t$. This derivation is conceptually similar to that of twist Jacobians Aoki et al. (2019); Li et al. (2021b) and the analytical warp regularization in NeRF Wang et al. (2023), effectively capturing continuous deformation over time.
In the context of Gaussian Splatting, the motion of Gaussians between two adjacent frames is bijective because all Gaussians are transformed from the same canonical space, creating a one-to-one correspondence among the Gaussians. Consequently, we can compute the velocity field directly from the forward warp field $\mathcal{W}_{c\rightarrow}(\mathcal{G},t)$. To analytically derive the velocity field $v(\mathcal{G};t)$, we must represent the gradient of the warp field in relation to time and warp parameters. Assuming that the Gaussian motions are small, we can utilize a modified Lucas-Kanade expression. By applying a Taylor series expansion to the warp field, we decompose the velocity into two essential components: the feature gradient of the warp field with respect to the corresponding Gaussian parameter and the analytical warp Jacobian Wang et al. (2023); Aoki et al. (2019). This relationship is expressed as:
$v(\mathcal{G};t)=\frac{d\mathcal{W}_{c\rightarrow}(\mathcal{G},t)}{dt}\Big{|}_{(\mathcal{G},t)}=J^{-1}_{\mathcal{G}}\frac{\partial\mathcal{W}_{c\rightarrow}(\mathcal{G};t)}{\partial t}.$ (10)
In this equation, the term $\frac{\partial\mathcal{W}_{c\rightarrow}(\mathcal{G};t)}{\partial t}$ indicates how each output Gaussian parameter varies with respect to the input canonical Gaussian parameters. Meanwhile, $J_{\mathcal{G}}$ describes how modifications in the warp field output influence the transformation of the Gaussians. In practical scenarios, the warp Jacobian can face numerical instabilities when $det(J_{\mathcal{G}})<\epsilon$, especially if the canonical Gaussian scene has not yet stabilized. Discarding these unstable motions can negatively impact subsequent densification and pruning processes. To mitigate these instabilities, we substitute the inverse Jacobian $J^{-1}_{\mathcal{G}}$ with the Moore–Penrose pseudoinverse $J^{+}_{\mathcal{G}}$.
The derived expression is valid for any Gaussian parameter input into the deformation network, provided the input and output dimensions are consistent. However, given that the scaling terms and spherical harmonics of the Gaussians show minimal changes between adjacent frames, we choose to omit these from the calculations for improved efficiency. Summarizing, we compute the velocity fields corresponding to the displacement and rotation of Gaussians as follows:
$v(\bm{\mu};t)=J^{+}_{\bm{\mu}}\frac{\partial\mathcal{W}_{c\rightarrow}(\bm{\mu};t)}{\partial t};\quad v(\bm{q};t)=J^{+}_{\bm{q}}\frac{\partial\mathcal{W}_{c\rightarrow}(\bm{q};t)}{\partial t},$ (11)
@@
where $v(\bm{\mu};t):\mathbb{R}^{3}\times\mathbb{R}\rightarrow\mathbb{R}^{3}$ and $v(\bm{q};t):\mathbb{R}^{4}\times\mathbb{R}\rightarrow\mathbb{R}^{4}$ describe the positional and rotational velocity of a Gaussian in the world coordinate, respectively.
@@ 4.2 Scene field from time integration
Time integration on the velocity field. Given the velocity field from equation 11, and 3D Gaussians $\mathcal{G}(t)$ observed at time step $t$, we can apply time integration using a Runge-Kutta Batschelet (1952) numerical solver on the velocity field to obtain the offset parameters for Gaussians at time $(t+\Delta t)$ Specifically, we employ Monte Carlo integration with a sampling size of 10 points over each domain, which provides an efficient and well-balanced alternative to deterministic approaches. The integration is formulated as:
@@
We then train the warp field network, identical to the one outlined in Yang et al. (2023), to minimize the difference between the analytically derived scene flow, and the ground truth scene flow decomposed into optical flow and depth describing the in-plane and out-of-plane Gaussian motions.
@@
For rendering the optical flow map $\hat{\mathcal{O}}_{t\rightarrow t+1}$, we want to calculate the motion’s influence on each pixel $\bm{u_{i}}$ from all Gaussians in the world coordinate. In the original 3D Gaussian Splatting, a pixel’s color is the weighted sum of the projected 2D Gaussian’s radiance contribution. Analogous to this formulation, the optical flow can be rendered by taking the weighted sum of the 2D Gaussians’ contribution to pixel shifts, as derived in previous literature on flow regularization Gao et al. (2024); Guo et al. (2024):
$\hat{\mathcal{O}}_{t\rightarrow t+1}=\Sigma_{i\in N}o_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})={\sum}_{i=1}w_{i}[\Sigma_{i,t_{2}}\Sigma^{-1}_{i,t_{1}}(\bm{u}_{t_{1}}-\bm{\mu}_{i,t_{1}})+\bm{\mu}_{i,t_{2}}-\bm{u}_{t_{1}}],$ (13)
where $\mathcal{O}_{i}$ denotes the optical flow of Gaussian $\mathcal{G}_{i}$ over the time interval, and $w_{i}=\frac{T_{i}\alpha_{i}}{\Sigma_{i}T_{i}\alpha_{i}}$ denotes the weighing factor of each Gaussian from $\alpha$-blending.
Similarly, the frame’s z-depth estimates $\hat{D}_{t+\Delta t}$ can be rendered from the discrete volume rendering approximation accounting for the per-Gaussian contributions:
$\hat{D}_{t+\Delta t}=\Sigma_{i\in N}d_{i,t_{2}}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (14)
where $d_{i,t_{2}}$ denotes the z-depth value of analytically derived Gaussian $\mathcal{G}_{i}$ from the viewing space at $(t+\Delta t)$. We normalize the optical flow and depth estimates for numerical stability. The optical flow and depth maps are rasterized simultaneously in the same forward pass with color. A performance comparison with various flow and depth priors can be found in A.3.
@@ 5 Experiments
In this section, we first provide the implementation details of the proposed warp field regulation and then validate our proposed method on five dynamic scene datasets captured with different levels of camera movements. Our method outperforms baseline approaches in both static and dynamic camera settings, achieving state-of-the-art results in quantitative and qualitative evaluations.
@@ 5.1 Implementation details
We implement the warp field using PyTorch Paszke et al. (2019), leveraging its Autodiff library for gradient and Jacobian computations. The framework is optimized with Adam Kingma & Ba (2015) as with 3DGS Kerbl et al. (2023) on a NVIDIA A100 Tensor Core Nvidia Corporation (2024). We use the Torchdiffeq Chen (2018) library for numerical integration. Modified from Deformable GS, the inputs to the warp field include 3D Gaussian positions ($\bm{\mu}$), time ($t$), and 3D Gaussian rotations ($\bm{q}$) in quaternions. Velocity derivations are enabled by removing the stop gradient operation from the network inputs. Positional encoding Mildenhall et al. (2021a) is applied to extend inputs’ frequency band Zheng et al. (2022). We refer readers to Appendix A.2 for more implementation details.
@@ 5.2 Evaluation
Benchmarked datasets. We evaluate our method using real-world monocular datasets with varying camera motion. These include DyCheck dataset Gao et al. (2022), captured with handheld monocular cameras; Dynamic Scene Yoon et al. (2020), captured by a stationary multi-view 8 camera rig with significant scene motion; Plenoptic Video Li et al. (2021c), captured using a static rig with 21 GoPro cameras GoPro (2023); Hypernerf Park et al. (2021b), which captures objects with moving topologies from a moving camera, suitable for quasi-static reconstruction; and sequences from DAVIS 2017 Perazzi et al. (2016) dataset containing near-static monocular videos.
Quantitative baseline comparisons. We evaluate our method’s novel-view synthesis capabilities on Plenoptic Videos, Hypernerf, Dynamic Scenes, and DyCheck, ordered by decreasing Effective Multi-view Factor (EMF) Gao et al. (2022). These datasets present increasing challenges, with DyCheck posing the greatest difficulty due to its low EMF and unreliable object motion. We first assess our method on Hypernerf and Plenoptics with higher EMFs against recent methods 3DGS Kerbl et al. (2023), Deformable GS Yang et al. (2023), 4DGaussians Wu et al. (2024), and MA-GS Guo et al. (2024). We report the standard metrics LPIPS Zhang et al. (2018), SSIM Wang et al. (2004), and PSNR Welstead (2009) in Table 6 and 4. We see that our method consistently outperforms these Gaussian baselines on scenes with higher EMFs.
@@
Table 2: Quantitative results on the DyCheck dataset.
Method Apple Spin Block Teddy Paper Windmill
NSFF 17.54 18.38 16.61 13.65 17.34
Marbles 17.57 15.49 16.88 13.57 18.67
4D Gaussians 15.41 14.41 11.28 12.36 15.60
Ours 16.03 16.71 15.46 13.60 17.41
Ours - pose refined 22.61 24.09 23.27 17.78 19.91
Table 3: Quantitative results for the Deformable GS method.
Method PSNR SSIM LPIPS
Deformable GS w/ flow 20.97 0.690 0.31
Deformable GS w/ depth 21.17 0.673 0.29
Deformable GS w/ both 23.94 0.702 0.28
Deformable GS 24.82 0.646 0.34
Ours 26.34 0.756 0.18
To further assess our method’s performance on monocular sequences with lower EMFs, we evaluate its performance against recent methods Marbles Stearns et al. (2024), NSFF Li et al. (2021d), NR-NeRF Tretschk et al. (2021b), Nerfies Park et al. (2021a), and Flow-supervised NeRF Wang_2023_CVPR, see Table 1 and 3. These comparisons highlight the effectiveness of our analytical scene flow approach compared to other flow-guided rendering frameworks. Notably, our method outperforms NSFF, which uses a similar supervision style but a different scene flow derivation. The comparison with Flow-supervised NeRF, which also uses analytical scene flow derived from an implicit radiance field, serves as an ablation study, demonstrating the effectiveness of our derived scene flow on an explicit scene representation. A.3.
@@
Table 6: Ablation study on scene flow decomposition on “Playground” scene from Dynamic Scenes dataset. See Sec. 5.2 for detailed descriptions of each ablated design.
@@
Comparisons with direct depth and flow Supervision. We note that it is essential to understand the performance improvements over regularizing the deformable network directly with depth and flow priors. We include ablation experiments with Deformable GS supervised with depth and flow losses, see Table 3. The results suggest that directly applying depth and flow regularizations encourages the scene to overfit to training viewpoints, leading to degraded performance. With depth regularizations, the scene geometries become more prominent, with a sacrifice in motion smoothness. Regularizing offsets between two frames can enforce accurate piecewise motion but still cannot generalize well to the discontinued unseen viewpoints. In Table 6, we ablate the analytical regularization and report the outcomes of the scene. As shown, each of the scene flow components helps in achieving high-quality synthesis. These ablation experiments reaffirm the validity of the proposed warp field regularization.
Trajectories from scene flow fields. We visualize the scene flow trajectories from Dynamic Scenes Dataset sequences to assess the quality and smoothness of the derived flow field. The Gaussians are subsampled from the dynamic regions of the canonical Gaussian space, and then time integrated to produce their respective displacements across time, visualized as colored trajectories in Fig. 3. We note that the trajectories are smooth and follow the expected motions of the dynamic objects. The visualized trajectories of the sequences demonstrate the framework’s potential to be extended for tracking in dynamic 3D Gaussian scenes. More trajectories on the DAVIS dataset will be shown in Appendix Fig. A7.
Geometric consistency. Fig. 4 compares our method with Deformable GS, which learns Gaussians’ time dependency without scene flow regularization. While Deformable GS renders visually accurate novel views, the underlying scene structures are inaccurate, leading to blurry dynamic objects and artifacts in depth maps. Our method, leveraging analytical scene flow regularization, achieves more accurate geometries, reflected in both visual and quantitative results (Table 1). This also enables accurate Gaussian motion tracking (Fig. 3). To explore alternative design choices, we compare with the direct flow and depth supervision on Deformable GS (Fig. 1). These comparisons indicate that our approach, which enforces regularization on the expected scene flow over a continuous time frame, results in more plausible scene structures compared to discrete time step supervision. We additionally conduct experiments on the DyCheck benchmark dataset, see Table 3. The additional visual comparison is shown in Appendix Fig. A2 with comparisons to the deformable baseline. Our results show that the analytical regularization significantly improves the render quality of the DyCheck dataset, especially when accurate camera poses are provided Wang et al. (2024).
Figure 5: Our method struggles with insufficient point cloud initialization due to reliance on non-rigid warping of scene geometry. In the “skating” scene, this results in unstable geometry at certain angles. Inaccurate camera calibrations also degrade our method, as shown in the “toby-sit” sequence, where miscalibration distorts the scene geometry.
@@ A.1 Scene Flow Regularization Formulation
Recall that in addition to the photometric regularization as detailed in the original 3D Gaussian Splatting Kerbl et al. (2023), $\mathscr{L}_{motion}$ is used as the structural regularization for the derived scene flow consisting of weighted depth and optical flow terms $\mathscr{L}_{motion}(t)=\alpha\mathscr{L}_{flow}+\beta\mathscr{L}_{depth}$.
The optical flow term minimizes the absolute error between the rendered optical flow map as computed from accounting for all Gaussians’ 2D projected motion contributions to each rendered pixel, and the reference flow from RAFT Teed & Deng (2020). In particular, the optical flow term is expressed as:
$\mathscr{L}_{flow}=|{\mathcal{O}}_{t\rightarrow t+1}-\hat{\mathcal{O}}_{t\rightarrow t+1}|.$ (15)
As noted in SparseNeRF Guangcong et al. (2023), monocular depth maps are scale-ambiguous thus directly using monocular depth priors with $p$-norm regularizations hinders the spatial coherency. We apply a local depth ranking loss on the rendered depth maps. We transfer the depth ranking knowledge from monocular depth $d_{mono}$ to rendered $d_{r}$ with:
$\mathscr{L}_{depth}=\sum_{d^{k_{1}}_{mono}<d^{k_{2}}_{mono}}max(d^{k_{1}}_{r}-d^{k_{2}}_{r}+m,0),$ (16)